# Monogamy of nonlocality from multipartite information causality

### v1

@@ Abstract
To find the maximum value of $\beta(\mathcal{B},\mathcal{E})$ given $\beta(\mathcal{A},\mathcal{B})$ permitted by the $\mathcal{IC}$ criteria (6),(7),(8) it suffices to consider tripartite no-signaling correlations $p(a,b,e|x,y,z)$ of the form,
@@
where $\alpha,\gamma\in[0,1]$ are convex coefficients of the PR-box correlations shared between $(\mathcal{A},\mathcal{B})$ and $(\mathcal{B},\mathcal{E})$, respectively, such that $\alpha+\gamma\leq 1$.
The proof follows from the data-processing inequality and has been deferred to the supplementary material for brevity. Notice that, for any point on the optimal slice specified by $\{\alpha,\gamma\}$ (9) the values CHSH functionals are specified as $\beta(\mathcal{A},\mathcal{B})=\frac{1+\alpha}{2}$ and $\beta(\mathcal{B},\mathcal{E})=\frac{1+\gamma}{2}$. Hence, our problem reduces to finding the maximum $\gamma$ given $\alpha$, such that the correlation (9) satisfies the $\mathcal{IC}$ criteria (6),(7),(8). These problems can now be efficiently tackled numerically, up to machine precision, as we describe below333codes containing the numerical solutions are available online at [33]. We plot the resultant curves in Figure 2. First, we discuss the case of the original (6) and the generalized (7) bipartite $\mathcal{IC}$ criteria.
Trivial monogamy from bipartite $\mathcal{IC}$:— Since the criteria (6) and (7) are essentially bipartite, a tripartite no-signaling correlation $p(a,b,e|x,y,z)$ must be locally post-processed into an effectively bipartite correlation $\tilde{p}(a^{\prime},b^{\prime}|x^{\prime},y^{\prime})$. Moreover, we need only consider deterministic post-processing schemes, referred to as wirings. A wiring is completely specified by the choice of the bipartition, for instance $\mathcal{A}$ and $\mathcal{B}^{\prime}\equiv(\mathcal{B},\mathcal{E})$, and the functions,
@@
where, $F_{i}:\{0,1\}^{n}\mapsto\{0,1\},\;\forall i\in\{1,2,3,4,5\}$. We note here that the grouped parties $\mathcal{B},\mathcal{E}$ may signal to each other. Consequently, a tripartite correlation $p(a,b,e|x,y,z)$ violates the bipartite criterion (7) for $\mathcal{IC}$, if there exists some wiring that produces an effectively bipartite correlation $\tilde{p}(a^{\prime},b^{\prime}|x^{\prime},y^{\prime})$ which violates it. In fact, such an approach has previously been used to study $\mathcal{IC}$ in multipartite Bell scenarios [14, 16, 32]. In particular, [18, 19, 20] claimed to have derived the quantum monogamy relation (5) through the bipartite $\mathcal{IC}$ criteria (6) by employing a wiring of the form (10). Contrary to these claims, we find that even the generalized $\mathcal{IC}$ criterion does not imply stronger-than-no-signaling monogamy relations, for any wiring of the form (10).
@@
Specifically, for all tripartite correlations $p(a,b,e|x,y,z)$ of the form (9) with $\alpha,\beta\in[0,1]$, we consider all possible wirings of the form (10), to retrieve the effectively bipartite correlations $\tilde{p}(a^{\prime},b^{\prime}|x^{\prime},y^{\prime})$. Then, for each such $\tilde{p}(a^{\prime},b^{\prime}|x^{\prime},y^{\prime})$, we employ the aforementioned protocol for the $(2\mapsto 1)$ RAC, with a binary symmetric noisy communication channel which flips the message bit $M$ with a probability $p(M^{\prime}=M\oplus 1|M)=\epsilon\in(1/2,1]$ for $M\in\{0,1\}$. Paralleling the observation in [15], we find that the tightest bounds on the maximum of value $\gamma$ given $\alpha\in[0,1]$ are recovered as $\epsilon\rightarrow 1/2$. We plot the resultant monogamy relation in Figure 2. The bipartite $\mathcal{IC}$ criteria (6),(7) retrieve the Tsirelson’s bounds, such that $\beta(\mathcal{B},\mathcal{E})\leq\beta_{Q}=\frac{1}{2}(1+\frac{1}{\sqrt{2}})$ for $\beta(\mathcal{A},\mathcal{B})\in[1/2,\frac{1}{2}(1-\frac{1}{\sqrt{2}})]$. However, for $\beta(\mathcal{A},\mathcal{B})\in[\frac{1}{2}(1-\frac{1}{\sqrt{2}}),\beta_{Q}]$ the monogamy relation implied by these criteria coincides with the no-signaling monogamy relation (4) such that, $\beta(\mathcal{B},\mathcal{E})\leq\frac{3}{2}-\beta(\mathcal{A},\mathcal{B})$. In particular, when $\mathcal{A},\mathcal{B}$ observe the Tsirelson’s bound $\beta(\mathcal{A},\mathcal{B})=\beta_{Q}$, the bipartite criteria (6),(7) fail to retrieve the quantum monogamy $\beta(\mathcal{B},\mathcal{E})\leq\frac{1}{2}$, as they allow for $\beta(\mathcal{B},\mathcal{E})\leq\frac{1}{2}(1-\frac{1}{\sqrt{2}})$. Thus, we conclude that the original (6) and generalized (7) bipartite criteria for $\mathcal{IC}$ fail to yield non-trivial monogamy relations beyond no-signaling. Let us now consider the multipartite $\mathcal{IC}$ criterion (8).
@@
Non-trivial monogamy from multipartite $\mathcal{IC}$:— For each tripartite correlation $p(a,b,e|x,y,z)$ of the form (9), we use the protocol for simplest $(N=3,n=2)$ multipartite communication task described above. Furthermore, we consider independent binary symmetric noisy classical channels between $\mathcal{A},\mathcal{B}$ and $\mathcal{E},\mathcal{B}$, respectively, which flip the input with probability $p(M^{\prime}_{1}=M_{1}\oplus 1|M_{1})=\epsilon_{1}$ and $p(M^{\prime}_{2}=M_{2}\oplus 1|M_{1})=\epsilon_{2}$, such that multipartite $\mathcal{IC}$ criterion (8) translates to,
@@
In contrast to the bipartite $\mathcal{IC}$ criteria (6),(7), we find that there exists correlations of the form (9), which satisfy the no-signalling monogamy relation (4), but violate the tripartite $\mathcal{IC}$ criterion (11) for some $\epsilon_{1},\epsilon_{2}\in(1/2,1]$. In other words, the tripartite $\mathcal{IC}$ criterion (11) implies a non-trivial monogamy relation of the form (1), which we plot in Figure 2. Most significantly, when $\mathcal{A},\mathcal{B}$ observe the Tsirelson’s bound $\beta(\mathcal{A},\mathcal{{B}})=\frac{1}{2}(1+\frac{1}{\sqrt{2}})$, the criterion (11) is violated for all $\beta(\mathcal{B},\mathcal{E})>1/2$, up to machine precision. This forms our first result,
@@
For $\beta(\mathcal{A},\mathcal{{B}})=\beta_{Q}$, the tripartite $\mathcal{IC}$ criterion (11) recovers the quantum monogamy (5) implying that $\mathcal{B}$ must be completely uncorrelated with the third party $\mathcal{E}$, such that, $\beta(\mathcal{B},\mathcal{E})=1/2$.
Moreover, the criterion (11) retrieves tighter than no-signaling monogamy relations for $\beta(\mathcal{A},\mathcal{B})\in[0.8333,\beta_{Q}]$. We now demonstrate that the monogamy relation implied by (11), although weaker than the quantum monogamy relation (5), is strong enough to enhance the information theoretic security of DIQKD protocols.
Secure DIQKD from $\mathcal{IC}$:— Recall that the generic monogamy relations of the form (2) can be used to derive the condition (3) for ensuring security of DIQKD protocol based on the CHSH scenario against individual attach of an eavesdropper constrained by a nonlocal theory $T$ [22, 6, 23, 24]. In Figure 2, we plot the condition (3) for ensuring security of DIQKD protocol. Observe that neither the no-signaling condition (4) nor the bipartite criteria (6),(7) yield security for realizable quantum correlations $\beta(\mathcal{A},\mathcal{B})\leq\beta_{Q}$. However, we find that the tripartite criterion (11) ensures secure DIQKD for a range of realizable quantum correlations, which forms our final result,
@@
For $\beta(\mathcal{A},\mathcal{B})\in[0.8471,\beta_{Q}]$, the monogamy relation implied by the tripartite $\mathcal{IC}$ criterion (11) satisfies the criterion (3), thereby, ensuring the security of DIQKD against no-signaling adversaries constrained by $\mathcal{IC}$.
Summarizing, we established formal connections between $\mathcal{IC}$ and monogamy of nonlocality. Specifically, we considered monogamy relations of the form (2) in the simplest tripartite Bell scenario. We used Lemma (1) to restrict to a two-parameter slice of the tripartite no-signaling polytope, on which we evaluated the $\mathcal{IC}$ criteria (6),(7),(8), to retrieve the implied monogamy relations (Figure 2). We remark here that since Lemma (1) holds for any informational criteria which satisfy the data-processing inequality, it opens the way for efficient evaluation of $\mathcal{IC}$ criteria beyond the ones considered here. We find that the bipartite criteria (6),(7) fail to retrieve a monogamy relation beyond no-signaling. Nevertheless, we demonstrate that the multipartite $\mathcal{IC}$ criteria (8),(11) can retrieve non-trivial monogamy relation stricter than the one implied by the no-signaling condition (4). Finally, we use this monogamy relation to demonstrate that the informational constraints implied by $\mathcal{IC}$ ensure the information theoretic security of realizable DIQKD protocols against individual attacks of an adversary.
Notably, the tripartite $\mathcal{IC}$ criterion (11) retrieves the quantum bound on $\beta({\mathcal{B},\mathcal{E}})$ when $\beta(\mathcal{A},\mathcal{B})=\beta_{Q}$. However, there remains a gap between the bounds implied by (11) and quantum monogamy (5), on $\beta_{\mathcal{B},\mathcal{E}}$ for the general case $\beta_{\mathcal{A},\mathcal{B}}\in(0.5,\beta_{Q})$. Owing to the inherent protocol-dependent formulation of the $\mathcal{IC}$ criteria, it remains an open question if this gap can be closed by employing different protocols, non-binary-symmetric noisy classical channels, or other $\mathcal{IC}$ criteria. Finally, it is interesting to explore the extension of our results to monogamy relations between other Bell inequalities in larger Bell scenarios.
Figure 2: A plots of the maximum value of the CHSH functional $\beta(\mathcal{B},\mathcal{E})$ implied by the monogamy relations (of the form (2)) considered in this work, against the CHSH functional $\beta(\mathcal{A},\mathcal{B})\in[1/2,1]$. The dashed and solid black lines represent the monogamy relations implied by the no-signaling condition (4) and quantum theory (5), respectively. The solid blue line represents monogamy relation implied by the bipartite $\mathcal{IC}$ criteria (6),(7) taking into account all possible wiring of the form (10). The solid orange line represents the monogamy relation implied by the tripartite $\mathcal{IC}$ criterion (11). Finally, the solid gray line exhibits the security condition (3) for DIQKD. Notice that, in contrast to the bipartite criterion, the tripartite $\mathcal{IC}$ criterion implies a non-trivial monogamy relation for $\beta(\mathcal{A},\mathcal{B})\in[0.8333,\beta_{Q}=\frac{1}{2}(1+\frac{1}{\sqrt{2}})$, and ensures security of DIQKD for $\beta(\mathcal{A},\mathcal{B})\in[0.8471,\beta_{Q}]$.
@@
We thank Rafael Rabelo for fruitful discussions and suggestions. This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001 - and by the Brazilian National Council for Scientific and Technological Development (CNPq). This work was partially supported by the Foundation for Polish Science (IRAP project, ICTQT, contract No. MAB/218/5, co-financed by EU within the Smart Growth Operational Programme). AC acknowledges financial support by NCN grant SONATINA 6 (contract No. UMO-2022/44/C/ST2/00081). C.R acknowledges support from the Narodowe Centrum Nauki (NCN) (SHENG project, contract No. 2018/30/Q/ST2/00625) and partially by the National Centre for Research and Development (QuantERA project, contract no. QUANTERA/2/2020). P.R.D acknowledges support from the NCN Poland, ChistEra-2023/05/Y/ST2/00005 under the project Modern Device Independent Cryptography (MoDIC).
@@
In this appendix, we provide proofs for the updated version of multipartite information causality ($\mathcal{IC}$) introduced in the equation (8) and Lemma 1. The main ingredient of this proof is an axiom of $\mathcal{IC}$, namely, the data processing inequality, which states that any local manipulation of data can only decay information, and is expressed as,
$I(A:B)\geq I(A:B^{\prime}),$ (12)
where $I(A:B)$ is an abstract theory independent mutual information between systems $A,B$, and $B^{\prime}$ is obtained from $B$ via a local transformation.
@@
As the $\mathcal{IC}$ criterion (7) generalizes the multipartite criterion in Eq. (8) in [17] by including the messages $\mathbf{M^{\prime}}\equiv\{M^{\prime}_{k}\}^{N-1}_{k=1}$, the proof closely follows the proof presented in Appendix B of [17]. Apart from the data-processing inqeuality (12) we also invoke the other two theory-independent axioms of $\mathcal{IC}$ in this proof, namely, consistency and the chain-rule presented in [12]. With the multipartite communication scenario depicted in Figure 1, let us consider the mutual information quantity $I(\mathbf{x}^{k}:\mathbf{x}^{1},\cdots,\mathbf{x}^{k-1},\mathbf{x}^{k+1},\cdots,\mathbf{x}^{N-1},\mathbf{M^{\prime}},c)$ which provides a measure of the entire network’s knowledge about the data set of part $k$. Using the relation provided in Eq. (B8) from [17], we derive the left-hand side of (8) in the following way,

### v2

@@ Abstract
To find the maximum CHSH value between $\mathcal{B}$ and $\mathcal{E}$, $\beta(\mathcal{B},\mathcal{E})$, permitted by information causality, when $\mathcal{A}$ and $\mathcal{B}$ witness a CHSH value, $\beta(\mathcal{A},\mathcal{B})$, it suffices to consider tripartite no-signaling correlations $p(a,b,e|x,y,z)$ of the form,
@@
where $\alpha,\gamma\in[0,1]$, and $\alpha+\gamma\leq 1$.
The proof follows from the data-processing inequality and has been deferred to the supplementary material for brevity. Notice that, for any point on the optimal slice specified by $\{\alpha,\gamma\}$ (9) the values CHSH functionals are specified as $\beta(\mathcal{A},\mathcal{B})=\frac{1+\alpha}{2}$ and $\beta(\mathcal{B},\mathcal{E})=\frac{1+\gamma}{2}$. Hence, our problem reduces to finding the maximum $\gamma$ given $\alpha$, such that the correlation (9) satisfies the $\mathcal{IC}$ criteria (6),(7),(8). These problems can now be efficiently tackled, up to machine precision, as we describe below 333codes containing the numerical solutions are available online at [39]. We plot the resultant curves in Figure 3.
Figure 2: Wiring procedure that takes tripartite correlations $p(a,b,e|x,y,z)$ and produces a bipartite effective one $p_{eff}(a^{\prime},b^{\prime}|x^{\prime},y^{\prime})$.
No monogamy from bipartite $\mathcal{IC}$:—Before presenting our main findings, we emphasize the necessity of a multipartite framework for $\mathcal{IC}$ in the context of monogamy relations. Specifically, we demonstrated in Figure 3 that both the original (6) and the generalized (7) bipartite formulations of $\mathcal{IC}$ fail to establish non-trivial monogamy relations for CHSH inequalities, such as (2). In this regard, we emphasize that the criteria (6) and (7) are essentially bipartite. Consequently, a tripartite no-signaling correlation $p(a,b,e|x,y,z)$ must be locally post-processed into an effectively bipartite correlation $\tilde{p}(a^{\prime},b^{\prime}|x^{\prime},y^{\prime})$ in order to be tested against them. Moreover, we need only consider deterministic post-processing schemes, referred to as wirings. A wiring is completely specified by the choice of the bipartition, for instance $\mathcal{A}$ and $\mathcal{B}^{\prime}\equiv(\mathcal{B},\mathcal{E})$, and the functions,
@@
where, $F_{i}:\{0,1\}^{n}\mapsto\{0,1\},\;\forall i\in\{1,2,3,4,5\}$. An example of such a wiring is depicted in Figure 2. We note here that the grouped parties $\mathcal{B},\mathcal{E}$ may signal to each other. Consequently, a tripartite correlation $p(a,b,e|x,y,z)$ violates the bipartite criterion (7) for $\mathcal{IC}$, if there exists some wiring that produces an effectively bipartite correlation $\tilde{p}(a^{\prime},b^{\prime}|x^{\prime},y^{\prime})$ which violates it. In fact, such an approach has previously been used to study $\mathcal{IC}$ in multipartite Bell scenarios [15, 17, 33]. In particular, [34, 35, 36] claimed to have derived the quantum monogamy relation (5) through the bipartite $\mathcal{IC}$ criteria (6) by employing a wiring of the form (10). Contrary to these claims, we find that neither the original (6) nor the generalized $\mathcal{IC}$ (7) criteria imply stronger-than-no-signaling monogamy relations, for any wiring of the form (10).
@@
Specifically, for all tripartite correlations $p(a,b,e|x,y,z)$ of the form (9) with $\alpha,\gamma\in[0,1]$, we consider all possible wirings of the form (10), to retrieve the effectively bipartite correlations $\tilde{p}(a^{\prime},b^{\prime}|x^{\prime},y^{\prime})$. Then, for each such $\tilde{p}(a^{\prime},b^{\prime}|x^{\prime},y^{\prime})$, we employ the aforementioned protocol for the $(2\mapsto 1)$ RAC, with a binary symmetric noisy communication channel which flips the message bit $M$ with a probability $p(M^{\prime}=M\oplus 1|M)=\epsilon\in[0,1/2)$ for $M\in\{0,1\}$. Paralleling the observation in [16], we find that the tightest bounds on the maximum of value $\gamma$ given $\alpha\in[0,1]$ are recovered as $\epsilon\rightarrow 1/2$. We plot the resultant monogamy relation in Figure 3, which is commented on in the main text. In this case, the bipartite $\mathcal{IC}$ criteria (6),(7) retrieve the Tsirelson’s bounds, such that $\beta(\mathcal{B},\mathcal{E})\leq\beta_{Q}=\frac{1}{2}(1+\frac{1}{\sqrt{2}})$ for $\beta(\mathcal{A},\mathcal{B})\in[1/2,\frac{1}{2}(2-\frac{1}{\sqrt{2}})]$. However, for $\beta(\mathcal{A},\mathcal{B})\in[\frac{1}{2}(2-\frac{1}{\sqrt{2}}),\beta_{Q}]$ the monogamy relation implied by these criteria coincides with the no-signaling monogamy relation (4) such that, $\beta(\mathcal{B},\mathcal{E})\leq\frac{3}{2}-\beta(\mathcal{A},\mathcal{B})$. Consequently, we conclude that the original (6) and generalized (7) bipartite criteria for $\mathcal{IC}$ fail to yield non-trivial monogamy relations beyond no-signaling. In particular, when $\mathcal{A},\mathcal{B}$ observe the Tsirelson’s bound $\beta(\mathcal{A},\mathcal{B})=\beta_{Q}$, the bipartite criteria (6),(7) play no role for monogamy relations. In this case, in fact, while quantum monogamy from (5) implies $\beta(\mathcal{B},\mathcal{E})\leq\frac{1}{2}$, the bipartite criteria yield $\beta(\mathcal{B},\mathcal{E})\leq\frac{1}{2}(2-\frac{1}{\sqrt{2}})$. As we shall demonstrate in this section, addressing this limitation requires the more refined approach for $\mathcal{IC}$ that inherently accounts for multipartite scenarios, as formalized in (8).
@@
Monogamy from multipartite $\mathcal{IC}$:— Building on this refined framework, we analyze tripartite correlations of the form (9) on a specific communication task. More precisely, for each tripartite correlation $p(a,b,e|x,y,z)$ of the form (9), we consider the protocol described above for the simplest non-trivial multipartite communication task with $(N=3,n=2)$. Furthermore, we assume independent binary symmetric noisy classical channels between $\mathcal{A},\mathcal{B}$ and $\mathcal{E},\mathcal{B}$, respectively, which flip the input with probability $p(M^{\prime}_{1}=M_{1}\oplus 1|M_{1})=\epsilon_{1}$ and $p(M^{\prime}_{2}=M_{2}\oplus 1|M_{1})=\epsilon_{2}$, such that multipartite $\mathcal{IC}$ criterion (8) translates to,
@@
Figure 3: A plot of the maximum value of the CHSH functional $\beta(\mathcal{B},\mathcal{E})$ implied by the monogamy relations (of the form (2)) considered in this work, against the CHSH functional $\beta(\mathcal{A},\mathcal{B})\in[1/2,1]$. The dashed and solid black lines represent the monogamy relations implied by the no-signaling condition (4) and quantum theory (5), respectively. The solid blue line represents monogamy relation implied by the bipartite $\mathcal{IC}$ criteria (6),(7) taking into account all possible wiring of the form (10). The solid orange line represents the monogamy relation implied by the tripartite $\mathcal{IC}$ criterion (11). Finally, the solid gray line exhibits the security condition (3) for DIQKD. Notice that, in contrast to the bipartite criterion, the tripartite $\mathcal{IC}$ criterion implies a non-trivial monogamy relation for $\beta(\mathcal{A},\mathcal{B})\in[0.8333,\beta_{Q}=\frac{1}{2}(1+\frac{1}{\sqrt{2}})]$, and ensures security of DIQKD for $\beta(\mathcal{A},\mathcal{B})\in[0.8471,\beta_{Q}]$.
In contrast to the bipartite $\mathcal{IC}$ criteria (6),(7), we find that there exists correlations of the form (9), which satisfy the no-signalling monogamy relation (4), but violate the tripartite $\mathcal{IC}$ criterion (11) for some $\epsilon_{1},\epsilon_{2}\in[0,1/2)$. In other words, the tripartite $\mathcal{IC}$ criterion (11) implies a non-trivial monogamy relation of the form (1), which we plot in Figure 3. Most significantly, when $\mathcal{A},\mathcal{B}$ observe the Tsirelson’s bound $\beta(\mathcal{A},\mathcal{{B}})=\frac{1}{2}(1+\frac{1}{\sqrt{2}})$, the criterion (11) is violated for all $\beta(\mathcal{B},\mathcal{E})>1/2$, up to machine precision. This forms our first result,
@@
When $\mathcal{A}$ and $\mathcal{B}$ witness the maximum quantum violation of the CHSH inequality such that $\beta(\mathcal{A},\mathcal{{B}})=\beta_{Q}$, tripartite information causality (11) implies $\mathcal{B}$ must be completely uncorrelated with any third party $\mathcal{E}$, such that the CHSH value between $\mathcal{B}$ and $\mathcal{E}$ must be $\beta(\mathcal{B},\mathcal{E})=1/2$, thereby recovering the quantum monogamy (5).
Moreover, the criterion (11) retrieves tighter than no-signaling monogamy relations for $\beta(\mathcal{A},\mathcal{B})\in[0.8333,\beta_{Q}]$. In the supplementary material, we show how the refinement provided in Eq. (8) plays a crucial role in the derivation of this result. We now demonstrate that the monogamy relation implied by (11), although weaker than the quantum monogamy relation (5), is strong enough to enhance the information-theoretic security of DIQKD protocols.
Secure DIQKD from $\mathcal{IC}$:— Recall that the generic monogamy relations of the form (2) can be used to derive the condition (3) for ensuring security of DIQKD protocol based on the CHSH scenario against individual attack of an eavesdropper constrained by a nonlocal theory $T$ [21, 6, 22, 23]. In Figure 3, we plot the condition (3) for ensuring security of DIQKD protocol. Observe that neither the no-signaling condition (4) nor the bipartite criteria (6),(7) yield security for realizable quantum correlations $\beta(\mathcal{A},\mathcal{B})\leq\beta_{Q}$. In contrast, we find that in the range $\beta(\mathcal{A},\mathcal{B})\in[0.8471,\beta_{Q}]$, the monogamy relation implied by the tripartite $\mathcal{IC}$ criterion (11) satisfies the criterion (3), thereby, ensuring the security of DIQKD, which forms our main result,
@@
Tripartite information causality (11) ensures security of DIQKD whenever $\mathcal{A}$ and $\mathcal{B}$ witness the CHSH value in the realizable range of quantum correlations, $\beta(\mathcal{A},\mathcal{B})\in[0.8471,\beta_{Q}]$.
The results establish a strong connection between a largely foundational field and practical applications. Notably, the security based on $\mathcal{IC}$ surpasses that based on no-signaling monogamy relations, and holds within the range $\beta(\mathcal{A},\mathcal{B})\in[0.8471,\beta_{Q}]$, which are attainable by means of quantum experiments. This raises the question of whether $\mathcal{IC}$-based security proofs can consistently outperform those grounded solely in the no-signaling principle. In this context, more refined security criteria and amplification protocols could significantly enhance these findings. Furthermore, since $\mathcal{IC}$ imposes tighter constraints than no-signaling, our results suggest a potential breakthrough in device-independent (DI) proofs within the quantum key distribution (QKD) paradigm.
To summarize, theoretical security proofs in device-independent quantum key distribution (DIQKD) are often based solely on the no-signaling principle without assuming any specific underlying physical theory describing the apparatus. However, strong evidence—such as that from the information causality ($\mathcal{IC}$) principle—suggests that a wide range of no-signaling correlations have no correspondence in nature, as they lead to implausible consequences. This observation raises the fundamental question of whether theoretical security proofs can also be drawn based on stronger-than-no-signaling reasonable principles. In consonance with this concern, here we demonstrated that when parties are assumed to be limited by the $\mathcal{IC}$ principle, security holds for a range of physically achievable parameters. A central open question regarding $\mathcal{IC}$ is whether it fully characterizes the set of quantum correlations. Nevertheless, our results reveal that, even in the negative case, if $\mathcal{IC}$ can be trusted as nature’s physical principle, secure cryptographic keys can be extracted against potential supra-quantum attacks. This follows directly from the nontrivial monogamy properties of Bell inequality violations derived from the multipartite formulation of $\mathcal{IC}$. Notably, the multipartite $\mathcal{IC}$ principle recovers the strong quantum monogamy of maximal CHSH inequality violations without relying on the Hilbert space formalism. Our findings highlight the necessity of the multipartite framework by demonstrating that bipartite formulations of $\mathcal{IC}$ alone do not impose monogamy constraints on CHSH inequalities. Given the inherent protocol-dependent to any formulation of the $\mathcal{IC}$ criteria, it remains open whether this gap can be closed through alternative protocols, as well as including non-binary-symmetric noisy classical channels in the analysis. In conclusion, building upon significant developments in the DIQKD literature regarding the no-signaling principle, our results provide a foundation for exploring the implications of stricter device-independent operational principles in practical applications, such as quantum key distribution (QKD) and quantum random number generation (QRNG). Promising natural directions for future research include analyzing security under general eavesdropping scenarios—such as collective or coherent attacks [19]—exploring amplification protocols [37], and extending our approach to more general cryptographic settings, including protocols based on chained Bell inequalities [38].
@@
We thank Rafael Rabelo for fruitful discussions and valuable suggestions. This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001 - and by the Brazilian National Council for Scientific and Technological Development (CNPq). This work is partially carried out under IRA Programme, project no. FENG.02.01-IP.05-0006/23, financed by the FENG program 2021-2027, Priority FENG.02, Measure FENG.02.01., with the support of the FNP. This work was partially supported by the Foundation for Polish Science (IRAP project, ICTQT, contract No. MAB/218/5, co-financed by EU within the Smart Growth Operational Programme). AC acknowledges financial support by NCN grant SONATINA 6 (contract No. UMO-2022/44/C/ST2/00081). C.R acknowledges support from the Narodowe Centrum Nauki (NCN) (SHENG project, contract No. 2018/30/Q/ST2/00625) and partially by the National Centre for Research and Development (QuantERA project, contract no. QUANTERA/2/2020). P.R.D acknowledges support from the NCN Poland, ChistEra-2023/05/Y/ST2/00005 under the project Modern Device Independent Cryptography (MoDIC).
@@
In this appendix, we provide proofs for the updated version of multipartite information causality ($\mathcal{IC}$) introduced in the equation (8), Lemma 1, and no-monogamy relations from bipartite formulated criteria for $\mathcal{IC}$.
@@
The main ingredients of this proof are the axioms of the original $\mathcal{IC}$ formulation, namely, the chain rule of the mutual information and the data processing inequality, which are respectively expressed as,
$I(A:B|C)=I(A:B,C)-I(A:C);$ (12)
$I(A:B^{\prime})\leq I(A:B),\quad\text{where}\quad B\longrightarrow B^{\prime}.$ (13)
The latter states that any local manipulation of data can only decay information where $I(A:B)$ is an abstract theory independent mutual information between systems $A,B$, and $B^{\prime}$ is obtained from $B$ via a local transformation. In addition to these two requirements, our proof also relies on another theory-independent axiom of $\mathcal{IC}$ (see [13]), known as consistency, which asserts that the mutual information $I(A:B)$ must reduce to Shannon’s mutual information when both subsystems $A$ and $B$ are classical. With the multipartite communication scenario depicted in Figure 1, let us consider the mutual information quantity of the bit-string of the $k$-th part, $\mathbf{x}^{k}$, with the data and remaining parties, local resources of the receiver, $c$, and all received messages $\{M^{\prime}_{k}\}^{N-1}_{k=1}$ denoted as $\mathbf{M^{\prime}}$, i.e., $I(\mathbf{x}^{k}:\mathbf{x}^{1},\cdots,\mathbf{x}^{k-1},\mathbf{x}^{k+1},\cdots,\mathbf{x}^{N-1},\mathbf{M^{\prime}},c)$. This quantity provides a measure of the entire network’s knowledge about the data set of part $k$. In this case, for brevity, we may consider the relation obtained in Ref. [18] (see Eq. (B8)), which states that when both chain rule and data processing conditions hold, we may straightly write,
@@
Notice that (14) is independent of the relation that the involved random variables may have among them.
@@
We stress that keeping this particular form for the multipartite $\mathcal{IC}$ criterion becomes crucial for the results presented in the main text. In fact, the previous formulation presented in [18] does not explicitly account for the received message on the measure of the receiver knowledge about the initial data (i.e., $I(X_{i}^{k}:X_{i}^{1},\dots,X_{i}^{k-1},X_{i}^{k+1},\dots,X_{i}^{N-1},G_{i})$). It becomes clear by analyzing the same scenario considered for the results of the main text in the noiseless case (three parties and the same protocol employed by the parties). In the extreme case where parties share a PR-box, $p(a,b,e|x,y,z)=\delta_{a\oplus b=xy}/4$, the receiver guessing might be written as $G_{i}=X_{i}^{1}\oplus M_{2}$, which leads the mutual information terms present in (8) to $I(X_{i}^{k}:X_{i}^{3-k},M_{1},M_{2},G_{i})=1$ , while for the previous criterion we will have $I(X_{i}^{k}:X_{i}^{3-k},G_{i})=0$. The same can be obtained considering the other extreme case where $p(a,b,e|x,y,z)=\delta_{e\oplus b=zy}/4$, and $G_{i}=X_{i}^{2}\oplus M_{1}$. Consequently, the previous multipartite criterion of [18] in terms of $I(X_{i}^{k}:X_{i}^{3-k},G_{i})$ witness no the violation of $\mathcal{IC}$ for the optimal slice (9), considering the same particular encoding and decoding protocols. Thus, being unable to recover the monogamy results presented in Fig.3.